I tried this but i think it's wrong can u please help again:
Sqrt x+9 +3=x
x+9 is in the radical(square root)

Answer by stanbon(75887)

Sqrt(x+9) +3=x
sqrt(x+9) = x-3
Square both sides to get:
x+9 = x^2-6x+9
x^2-7x=0
x(x-7)=0
x=0 or x=7
Check answer in the orignal equation:
If x=0, sqrt(0+9)=0-3
Wrong as sqrt9 = +3
---------
If x=7, sqrt(7+9)=7-3
sqrt16=4
4=4
Good solution
-----------
Final solution:
x=7
Cheers,
Stan H.